OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? x - y = 8 -x - y = -7

OpenStudy (anonymous):

x=8+y

OpenStudy (ash2326):

@Suzan987654321 We have the equation \[x-y=8\] Add both sides by y. Can you find what's x?

OpenStudy (anonymous):

x = 8 + y

OpenStudy (anonymous):

???

OpenStudy (ash2326):

Great:D

OpenStudy (anonymous):

:) thank u !

OpenStudy (ash2326):

You did understand how to do, didn't you?

OpenStudy (anonymous):

ya i did